Let u = 4i-6 j and v=−8i+4j. Find u−v.
Answer #1

u - v = (4i - 6j) - (-8i + 4j)

u - v = (4 + 8)i + (-6 - 4)j = 12i - 10j

Comment in case any doubt please rate my answer....